# Modelling Non-monotone Risk Aversion and Convex Compensation in Incomplete Markets

### v1

@@ Abstract
In hedge funds, convex compensation schemes are popular to stimulate a high-profit performance for portfolio managers. In economics, non-monotone risk aversion is proposed to argue that individuals may not be risk-averse when the wealth level is low. Combining these two ingredients, we study the optimal control strategy of the manager in incomplete markets. Generally, we propose a wide class of utility functions, the Piecewise Symmetric Asymptotic Hyperbolic Absolute Risk Aversion (PSAHARA) utility, to model the two ingredients, containing both non-concavity and non-differentiability as some abnormalities. Significantly, we derive an explicit optimal control for the family of PSAHARA utilities. This control is expressed into a unified four-term structure, featuring the asymptotic Merton term and the risk adjustment term. Furthermore, we provide a detailed asymptotic analysis and numerical illustration of the optimal portfolio. We obtain the following key insights: (i) A manager with the PSAHARA utility becomes extremely risk-seeking when his/her wealth level tends to zero; (ii) The optimal investment ratio tends to the Merton constant as the wealth level approaches infinity and the negative Merton constant when the wealth falls to negative infinity, implying that such a manager takes a risk-seeking investment as the wealth falls negatively low; (iii) The convex compensation still induces a great risk-taking behavior in the case that the preference is modeled by SAHARA utility. Finally, we conduct a real-data analysis of the U.S. stock market under the above model and conclude that the PSAHARA portfolio is very risk-seeking and leads to a high return and a high volatility (two-peak Sharpe ratio).
Keywords: Piecewise symmetric asymptotic hyperbolic absolute risk aversion (PSAHARA) utility, Optimal control, Asymptotic analysis, Empirical financial analysis
@@ 1 Introduction
Compensation incentive schemes are commonly adopted to share profits between the portfolio manager and the investors in hedge funds. A typical setting in the financial industry is the “2–20” scheme—$2\%$ management fee of the fund value and $20\%$ performance fee of the excess profit, where the latter usually takes the form of a call option; see Carpenter (2000). Such a compensation scheme gives the manager a strong incentive to chase a good performance. Mathematically, this scheme is a piecewise linear convex function and hence is referred to as a convex compensation scheme (formally in Eq. (5)). Multiple studies have explored the impact of incentive options on the investment strategies of fund managers; see Berkelaar, Kouwenberg and Post (2004), Hodder and Jackwerth (2007), Bichuch and Sturm (2014), He and Kou (2018) and Liang and Liu (2020). In these studies, fund managers are typically assumed to have constant relative risk aversion (CRRA) or hyperbolic absolute risk aversion (HARA) types of utilities or S-shaped utilities with piece power function variants.
@@
If the portfolio manager has a SAHARA preference, the actual utility is the composition of the SAHARA preference and the convex compensation scheme. This actual utility may not be of SAHARA on the whole domain and become complicated. To model this actual utility and further enhance the generality, we introduce a wide class of piecewise SAHARA (PSAHARA) utility functions. Roughly speaking, a utility function in the PSAHARA family satisfies that there is a partition of the real line making the utility SAHARA on each interval (while the different sections allow different parameter settings). Starting from this point, our contribution is fourfold:
First, we design a unified parameterization and introduce the PSAHARA utility family (Definition 2). We demonstrate that a PSAHARA utility inherits the non-monotone risk aversion characteristics from the SAHARA utility, allowing for studies on more flexible wealth levels and more complicated risk-taking behaviors. In the PSAHARA family, we emphasize the role of various threshold wealth levels and argue that the individual becomes less risk-averse below each threshold. By introducing some properties on the transformation-invariance of the PSAHARA utility, this wide family incorporates the above objective (composition of the SAHARA preference and the piecewise linear convex compensation) in hedge fund management as a motivating example (Propositions 12). The utility family may promote other application contexts of the utility theory.
Second, we formulate the continuous-time portfolio selection problem in incomplete Black–Scholes markets and obtain an explicit optimal control for the general PSAHARA utility (Theorems 12). As suggested above, the PSAHARA utility may not be of SAHARA on the whole domain, and may not necessarily be concave or differentiable. This creates difficulties in solving an explicit form of the portfolio. Technically in Proposition 3, we apply and extend the martingale and duality method (Karatzas, Lehoczky, Shreve and Xu (1991)) and the concavification techniques (Carpenter (2000)) to the incomplete markets to obtain the optimal portfolio. The optimal control is expressed in a “partial” feedback form, including parts of the optimal wealth and a unique Lagrange multiplier ($y^{*}$ in later contexts). We name the four components of the optimal portfolio by their mathematical and economic implications: the asymptotic Merton term, the risk adjustment term (due to the scale parameter in PSAHARA utility), the first-order risk aversion term (due to non-differentiability), and the loss aversion term (due to threshold wealth levels).
@@
Fourth, we conduct an empirical study on the motivating example in hedge fund management. We first give the corresponding optimal portfolio (Theorem 4). We apply this portfolio strategy in the real data of the U.S. stock market, using different estimation methods of the volatility process $\{\boldsymbol{\sigma}_{t}\}_{0\leqslant t\leqslant T}$. We analyze the Sharpe ratio of investment and find that there exists a “gambling” behavior on the linear segments of the composed utility. This means that even if the manager has a SAHARA utility, the convex compensation indeed induces a great risk-taking behavior (which coincides with the classic result of Carpenter (2000)). We further find a two-peak pattern of the corresponding Sharpe ratio, meaning that the PSAHARA portfolio leads to a high return and a high volatility.
We compare our novelties with the literature. In the study of incomplete markets, Pagés (1987) pioneers the study on optimal investment and adopts the martingale analysis approach. This approach is developed by He and Pearson (1991) and Karatzas, Lehoczky, Shreve and Xu (1991). These works mainly focus on the existence of optimal investment strategies, while we technically get rid of the assumption of the positive domain in Proposition 3 and provide an explicit optimal portfolio in incomplete markets in Theorem 2. In addition, Liang, Liu, Ma and Vinoth (2024) proposes a family of utility functions, “piecewise HARA (PHARA) utility”, which means that the utility is of HARA in each part of the domain. Our PSAHARA utility family shares a similar logic of generalization, but the SAHARA utility has a more complicated expression and the portfolio is more subtle to obtain. Furthermore, as the SAHARA utility can (asymptotically) reduce to a HARA utility, our PSAHARA utility contains the big family of the PHARA utility. Technically, the corresponding optimal control formula in Theorem 2 is designed in a way that can exactly reduce to the optimal control of the PHARA portfolio in Theorem 1 of Liang, Liu, Ma and Vinoth (2024). This makes the portfolio formula consistent and unified to implement in future research and financial practice.
This paper is structured as follows. In Section 2, we introduce our motivating example. We give the definitions of the SAHARA utility and incentive contracts. The model settings are introduced in Section 3. In Section 4, we present the explicit formula of the optimal portfolio, which is the main theorem of our study. We conduct the asymptotic analysis in Section 5, followed by the numerical illustration in Section 6. Finally, we revisit our motivating example and conduct empirical studies on the corresponding optimal portfolio in Section 7. All the proofs are included in Appendix A. Methods of the empirical study are in Appendix B.
@@ 2 Motivating Example in Hedge Funds
In this section, we propose a motivating example and give an analytical model of non-monotone risk aversion and convex compensation. The non-monotone risk aversion is depicted by the SAHARA utility class (Chen, Pelsser and Vellekoop (2011)), which has the following definition.
@@
A utility function $U$ with domain $\mathbb{R}$ is of the SAHARA class if its absolute risk aversion function $\text{ARA}\left(x\right)=-U^{{}^{\prime\prime}}\left(x\right)/U^{{}^{\prime}}\left(x\right)$ is defined on $\mathbb{R}$ and satisfies
$\text{ARA}\left(x\right)=\frac{\alpha}{\sqrt{\beta^{2}+\left(x-d\right)^{2}}}\geqslant 0,$ (1)
for given $\alpha\geqslant 0$ (the risk aversion parameter), $\beta>0$ (the scale parameter), and $d\in\mathbb{R}$ (the threshold wealth). To give the explicit expression of such a utility, we derive that there exist constants $c_{1}\in\mathbb{R}$, $c_{2}>0$ such that $U\left(x\right)=c_{1}+c_{2}\hat{U}\left(x\right)$ with
$\hat{U}(x;\alpha,\beta,d)=\left\{\begin{aligned} &-\frac{1}{\alpha^{2}-1}\left(\left(x-d\right)+\sqrt{\beta^{2}+\left(x-d\right)^{2}}\right)^{-\alpha}\left(\left(x-d\right)+\alpha\sqrt{\beta^{2}+\left(x-d\right)^{2}}\right),\quad\alpha\neq 1;\\ &\frac{1}{2}\log\left(\left(x-d\right)+\sqrt{\beta^{2}+\left(x-d\right)^{2}}\right)+\frac{1}{2}\beta^{-2}\left(x-d\right)\left(\sqrt{\beta^{2}+\left(x-d\right)^{2}}-\left(x-d\right)\right),\quad\alpha=1.\end{aligned}\right.$ (2)
@@
which reduces to a HARA utility on the interval of $(d,\infty)$. For better analytical tractability, in the case $\beta=0$, we let $x\in(d,\infty)$ be the domain of the utility function; see, e.g., Remark 7.
Current studies on the SAHARA utility (Chen, Pelsser and Vellekoop (2011) and others) usually assume $d=0$ for notation simplicity. However, we emphasize the role of the parameter $d$ as it represents different threshold wealth levels; see later Definition 2. As a result, we can have more general properties of the SAHARA class.
@@
$\Theta\left(x\right)=w\left(x-B_{T}\right)^{+}+vx,$ (5)
where $B_{T}$ is the discounted benchmark level, $w>0$ denotes the number of options and $10<v<$ 1 is the rate of management fee. In practice, we usually have that $10<v<w<$ 1 since $w$ represents the incentive compensation while $v$ represents the regular management fee. This is similar to the model in Hodder and Jackwerth (2007) but excludes the lower-bound liquidation boundary. Let $X_{T}$ denote the terminal fund value. Then the manager’s wealth is $\Theta(X_{T})$. We assume that the manager has a SAHARA utility $\hat{U}$. Hence, the utility function of the manager with respect to the terminal value of the fund under the compensation scheme is given by
$U\left(X_{T}\right):=\hat{U}\circ\Theta\left(X_{T}\right)=\left\{\ \begin{aligned} &\hat{U}\left(vX_{T}\right),&&x\leqslant B_{T};\\ &\hat{U}\left(w\left(X_{T}-B_{T}\right)+vX_{T}\right),&&x>B_{T},\end{aligned}\right.$ (6)
@@
It is a regular approach to set a lower bound or a liquidation boundary ($X_{T}>0$ a.s.) in the literature; see e.g. Carpenter (2000) and Bichuch and Sturm (2014). They assume the Inada condition on their utility function $U$. That is, $U$ is defined on $(0,\infty)$ and
@@
By shifting on the $x$-axis, Eq. (7) can be made valid for any real number. This assumption, though providing analytical tractability in their setting, excludes the cases where the fund value tends extremely low. Here, we do not set such a lower bound. Our setting includes the whole real line as the domain of the wealth level. We can hence deal with more comprehensive scenarios.
The explicit form is
@@
which belongs to the PSAHARA class introduced later in Section 3. This is further illustrated in Figure 1.
@@
We can further delve into the risk aversion behaviors of the manager:
$\text{ARA}\left(x\right)=\left\{\begin{aligned} &\frac{v\alpha}{\sqrt{\beta^{2}+\left(x-d\right)^{2}}},&&X_{T}\leqslant B_{T};\\ &\frac{\left(w+v\right)\alpha}{\sqrt{\beta^{2}+\left(x-wB_{T}-d\right)^{2}}},&&X_{T}>B_{T}.\end{aligned}\right.$ (9)
@@ 3.1 Market Model
We consider a multi-dimensional Black–Scholes model. To solve the explicit solution, all the market parameters are deterministic processes. There are a risk-free asset and risky assets in the market. The risk-free asset has a deterministic return rate and no volatility, while the risky assets have higher expected return rates and positive volatilities. We denote by the filtered probability space $\left(\Omega,\mathcal{F}_{T},\{\mathcal{F}_{t}\}_{0\leqslant t\leqslant T},\mathbb{P}\right)$ the financial market. The filtration $\{\mathcal{F}_{t}\}_{0\leqslant t\leqslant T}$ is the one generated by a $q$-dimensional standard independent Brownian motion $\{\mathbf{W}_{t}\}_{0\leqslant t\leqslant T}=\{\left(W_{1,t},\ldots,W_{d,t}\right)^{\intercal}\}_{0\leqslant t\leqslant T}$ and further augmented by all $\mathbb{P}$-null sets. Let the deterministic process $\{r_{t}\}_{0\leqslant t\leqslant T}$ denote the risk-free rate. The risk-free asset $\{S_{0,t}\}_{0\leqslant t\leqslant T}$ satisfies
@@
In the traditional incomplete Black–Scholes model, the market price of risk $\{\boldsymbol{\tilde{\theta}}_{t}\}_{0\leqslant t\leqslant T}$ can be any solution of the linear system:
$\boldsymbol{\sigma}_{t}\boldsymbol{\tilde{\theta}}_{t}=\boldsymbol{\mu}_{t}-r_{t}\mathbf{1}_{m}.$ (13)
Hence, if $m<q$, $\boldsymbol{\tilde{\theta}}_{t}$ is not unique. We choose Eq. (12), one of the solutions of Eq. (13), to represent the price of risk. Later we will show that this simplification does not influence the validity of our main theorems.
Next, we define the pricing kernel process $\{\xi_{t}\}_{0\leqslant t\leqslant T}$ as follows:
$\xi_{t}:=\exp\left\{-\int_{0}^{t}\left(r_{s}+\frac{1}{2}\|\boldsymbol{\theta}_{s}\|_{2}^{2}\right)\,\mathrm{d}s-\int_{0}^{t}\boldsymbol{\theta}_{s}^{\intercal}\,\mathrm{d}\mathbf{W}_{s}\right\},\quad 0\leqslant t\leqslant T,$ (14)
where $\|\boldsymbol{\theta}_{t}\|_{2}:=\left(\sum_{i=1}^{q}\theta_{i,t}\right)^{\frac{1}{2}}$. We denote by $\{\pi_{i,t}\}_{0\leqslant t\leqslant T}$ the amount of money invested in $i$th risky asset $S_{i}$ at time $t$. The wealth process $\{X_{t}\}_{0\leqslant t\leqslant T}$ is uniquely determined by the investment process $\{\boldsymbol{\pi}_{t}=\left(\pi_{1,t},\ldots,\pi_{m,t}\right)\}_{0\leqslant t\leqslant T}$ and an initial value $x_{0}>0$:
@@
A portfolio $\boldsymbol{\pi}=\{\boldsymbol{\pi}_{t}\}_{0\leqslant t\leqslant T}$ is called admissible if and only if $\boldsymbol{\pi}$ is an $\{\mathcal{F}_{t}\}_{0\leqslant t\leqslant T}$progressively measurable $\mathbb{R}^{m}$-valued process and $\int_{0}^{T}\|\boldsymbol{\pi}_{t}\|_{2}^{2}\,\mathrm{d}t<\infty$ almost surely. The decision maker conducts portfolio selection by solving the expected utility maximization problem:
@@
where $U$ is the composed utility function and $\mathcal{V}$ is the collection of all admissible portfolios.

### v2

# PSAHARA Utility Family: Modeling Non-monotone Risk Aversion and Convex Compensation in Incomplete Markets

@@ Abstract
In hedge funds, convex compensation schemes are adopted to stimulate a high-profit performance for portfolio managers. In economics, non-monotone risk aversion is proposed to argue that individuals may not be risk-averse when the wealth level is low. Combining these two ingredients, we study the optimal control strategy of the manager in incomplete markets. Generally, we propose a wide family of utility functions, the piecewise symmetric asymptotic hyperbolic absolute risk aversion (PSAHARA) utility, to model the two ingredients, containing both non-concavity and non-differentiability as some abnormalities. Technically, we propose an additional assumption and prove concavification techniques of non–concave utility functions with a left unbounded domain in incomplete markets. Next, we derive an explicit optimal control for the family of PSAHARA utilities. This control is expressed into a unified four-term structure, featuring the asymptotic Merton term. Furthermore, we provide a detailed asymptotic analysis and numerical illustration of the optimal portfolio. We obtain several key insights, including that the convex compensation still induces a great risk-taking behavior in the case that the preference is modeled by SAHARA utility. Finally, we conduct a real-data analysis of the U.S. stock market under the above model and conclude that the PSAHARA portfolio is very risk-seeking and leads to a high return and a high volatility.
Keywords: Utility theory, Piecewise symmetric asymptotic hyperbolic absolute risk aversion (PSAHARA) utility, Concavification technique, Asymptotic analysis, Empirical financial analysis
@@ 1 Introduction
Compensation incentive schemes are commonly adopted to share profits between the portfolio manager and the investors in hedge funds. A typical setting in the financial industry is the “2–20” scheme—$2\%$ management fee of the fund value and $20\%$ performance fee of the excess profit, where the latter usually takes the form of a call option; see Carpenter (2000). Such a compensation scheme gives the manager a strong incentive to chase a good performance. Mathematically, this scheme is a piecewise linear convex function and hence is referred to as a convex compensation scheme (formally in Eq. (5)). Multiple studies have explored the impact of incentive options on the investment strategies of fund managers; see Berkelaar, Kouwenberg and Post (2004), Hodder and Jackwerth (2007), Bichuch and Sturm (2014), He and Kou (2018) and Liang and Liu (2020). In these studies, fund managers are typically assumed to have constant relative risk aversion (CRRA) or hyperbolic absolute risk aversion (HARA) types of utilities or S-shaped utilities with piecewise power function variants.
@@
If the portfolio manager has a SAHARA preference, the actual utility is the composition of the SAHARA preference and the convex compensation scheme. This actual utility may not be of SAHARA on the whole domain and become complicated. To model this actual utility and further enhance the generality, we introduce a wide family of piecewise SAHARA (PSAHARA) utility functions. Roughly speaking, a utility function in the PSAHARA family satisfies that there is a partition of the real line making the utility SAHARA on each interval (while the different sections allow different parameter settings). Starting from this point, our contribution is fourfold.
First, we design a unified parameterization and introduce the PSAHARA utility family (Definition 2). Current studies on continuous-time portfolio selection adopt various criterion to evaluate risk and return, e.g., the risk measure in He, Jin and Zhou (2015), the S-shape utility in Dong and Zheng (2020), and the performance ratio in Guan, Liang and Xia (2023). Here we capture the individual preference by the PSAHARA utility family, which distinguishes from the commonly used CRRA, CARA, and HARA utilities by the non-monotone risk aversion characteristics inherited from the SAHARA utility, allowing for studies on more flexible wealth levels and more complicated risk-taking behaviors. In the PSAHARA family, we emphasize the role of various threshold wealth levels and argue that the individual becomes less risk-averse below each threshold. By introducing some properties on the transformation-invariance of the PSAHARA utility, this wide family incorporates the above objective (composition of the SAHARA preference and the piecewise linear convex compensation) in hedge fund management as a motivating example (Propositions 12). The utility family may promote other application contexts of the utility theory.
Second, we formulate the continuous-time portfolio selection problem in incomplete Black–Scholes markets and obtain an explicit optimal control for the general PSAHARA utility (Theorems 12). As suggested above, the PSAHARA utility may not be of SAHARA on the whole domain, and may not necessarily be concave or differentiable. This creates difficulties in solving an explicit formula of the portfolio. In the proofs, we explain these issues in details. Technically, in Theorem 1 and Proposition 5, we apply and extend the martingale and duality method (Karatzas, Lehoczky, Shreve and Xu (1991)) and rigorously discuss the concavification techniques (Carpenter (2000)) in the incomplete markets to obtain the optimal portfolio. To rigorously apply the martingale and duality method, we propose an additional assumption (Assumption 2), which guarantees the square integrability of the portfolio control processes; see Section 4.2 for details. The optimal control is expressed in a “partial” feedback form, including parts of the optimal wealth and a unique Lagrange multiplier ($y^{*}$ in later contexts). We name the four components of the optimal portfolio by their mathematical and economic implications: the asymptotic Merton term, the risk adjustment term (due to the scale parameter in PSAHARA utility), the first-order risk aversion term (due to non-differentiability), and the loss aversion term (due to threshold wealth levels).
@@
Fourth, we conduct an empirical study on the motivating example in hedge fund management. We first give the corresponding optimal portfolio (Theorem 4). We apply this portfolio strategy with the real data of the U.S. stock market, using different estimation methods of the volatility process $\{\boldsymbol{\sigma}_{t}\}_{0\leqslant t\leqslant T}$. We analyze the Sharpe ratio of investment and find that there exists a “gambling” behavior on the linear segments of the composed utility. This means that even if the manager has a SAHARA utility, the convex compensation indeed induces a great risk-taking behavior (which coincides with the classic result of Carpenter (2000)). We further find a two-peak pattern of the corresponding Sharpe ratio, meaning that the PSAHARA portfolio leads to a high return and a high volatility.
We compare our novelties with the literature. In the study of incomplete markets, Pagés (1987) pioneers the study on optimal investment and adopts the martingale analysis approach. This approach is developed by He and Pearson (1991) and Karatzas, Lehoczky, Shreve and Xu (1991). These works mainly focus on the existence of optimal investment strategies, while we technically get rid of the assumption of the left bounded domain of non–concave utility functions in Proposition 5 and provide an explicit optimal portfolio in incomplete markets in Theorem 2. In addition, Liang, Liu, Ma and Vinoth (2024) proposes a family of utility functions, “piecewise HARA (PHARA) utility”, which means that the utility is of HARA on each part of the domain. Our PSAHARA utility family shares a similar logic of generalization, but the SAHARA utility has a more complicated expression and the portfolio is more subtle to obtain. Furthermore, as the SAHARA utility can (asymptotically) reduce to a HARA utility, our PSAHARA utility contains the big family of the PHARA utility. Technically, the corresponding optimal control formula in Theorem 2 can exactly reduce to the optimal control of the PHARA portfolio in Theorem 1 of Liang, Liu, Ma and Vinoth (2024). This makes the portfolio formula consistent and unified to implement in future research and financial practice.
This paper is structured as follows. In Section 2, we introduce our motivating example. We give the definitions of the SAHARA utility and incentive contracts. The model settings are introduced in Section 3. In Section 4.1, we present the explicit formula of the optimal portfolio, which is the main theorem of our study. Section 4.2 shows our main technical contribution. We conduct the asymptotic and numerical analysis in Section 5. Finally, we revisit our motivating example and conduct empirical studies on the corresponding optimal portfolio in Section 6. All the proofs are included in Appendix A. Methods of the empirical study are in Appendix B.
@@ 2 Motivating Example in Hedge Funds
In this section, we propose a motivating example and give an analytical model of non-monotone risk aversion and convex compensation. The non-monotone risk aversion is depicted by the SAHARA utility family (Chen, Pelsser and Vellekoop (2011)). It characterizes complex risk attitudes and show distinct features compared to the widely studied CRRA, CARA and HARA utilities.
@@
A utility function $U$ with the domain $\mathbb{R}$ is of the SAHARA family if its absolute risk aversion function $\text{ARA}\left(x\right)=-U^{{}^{\prime\prime}}\left(x\right)/U^{{}^{\prime}}\left(x\right)$ is defined on $\mathbb{R}$ and satisfies
$\text{ARA}\left(x\right)=\frac{\alpha}{\sqrt{\beta^{2}+\left(x-d\right)^{2}}}\geqslant 0,\quad x\in\mathbb{R},$ (1)
for given $\alpha\geqslant 0$ (the risk aversion parameter), $\beta>0$ (the scale parameter), and $d\in\mathbb{R}$ (the threshold wealth). To give the explicit expression of such a utility, we derive that there exist constants $c_{1}\in\mathbb{R}$ and $c_{2}>0$ such that $U\left(x\right)=c_{1}+c_{2}\hat{U}\left(x\right)$ with
$\hat{U}(x;\alpha,\beta,d)=\left\{\begin{aligned} &-\frac{1}{\alpha^{2}-1}\left(\left(x-d\right)+\sqrt{\beta^{2}+\left(x-d\right)^{2}}\right)^{-\alpha}\left(\left(x-d\right)+\alpha\sqrt{\beta^{2}+\left(x-d\right)^{2}}\right),\quad\alpha\neq 1;\\ &\frac{1}{2}\log\left(\left(x-d\right)+\sqrt{\beta^{2}+\left(x-d\right)^{2}}\right)+\frac{1}{2}\beta^{-2}\left(x-d\right)\left(\sqrt{\beta^{2}+\left(x-d\right)^{2}}-\left(x-d\right)\right),\quad\alpha=1,\end{aligned}\right.$ (2)
@@
which reduces to a HARA utility on the interval of $(d,\infty)$. For better analytical tractability, in the case $\beta=0$, we let $x\in(d,\infty)$ be the domain of the utility function; see also Assumption 3 and Remark 8.
Current studies on the SAHARA utility (Chen, Pelsser and Vellekoop (2011) and others) usually assume $d=0$ for notation simplicity. However, we emphasize the role of the parameter $d$ as it represents different threshold wealth levels; see later Definition 2. As a result, we can have more general properties of the SAHARA family.
@@
$\Theta\left(x\right)=w\left(x-B_{T}\right)^{+}+vx,\quad x\in\mathbb{R},$ (5)
where $B_{T}$ is the discounted benchmark level, $w>0$ denotes the number of options and $10<v<$ 1 is the rate of management fee. In practice, we usually have that $10<v<w<$ 1 since $w$ represents the incentive compensation while $v$ represents the regular management fee. This is similar to the model in Hodder and Jackwerth (2007) and Chen, Hieber and Nguyen (2019) but excludes the lower-bound liquidation boundary. Let $X_{T}$ denote the terminal fund value. Then the manager’s wealth is $\Theta(X_{T})$. We assume that the manager has a SAHARA utility $\hat{U}$. Hence, the utility function of the manager with respect to the terminal value of the fund under the compensation scheme is given by
$U\left(X_{T}\right):=\hat{U}\circ\Theta\left(X_{T}\right)=\left\{\ \begin{aligned} &\hat{U}\left(vX_{T}\right),&&X_{T}\leqslant B_{T};\\ &\hat{U}\left(w\left(X_{T}-B_{T}\right)+vX_{T}\right),&&X_{T}>B_{T},\end{aligned}\right.$ (6)
@@
It is a regular approach to set a lower bound or a liquidation boundary ($X_{T}\geqslant 0$ a.s.) in the literature; see, e.g., Carpenter (2000) and Bichuch and Sturm (2014) and later Assumption 3. They assume the Inada condition on their utility function $U$. That is, $U$ is defined on $(0,\infty)$ and
@@
By shifting on the $x$-axis, Eq. (7) can be made valid for any real number. This assumption, though providing analytical tractability, excludes the cases where the fund value tends extremely low. Here, we do not set such a lower bound. Our setting includes the whole real line as the domain of the wealth level. We can hence deal with more comprehensive scenarios.
The explicit form of Eq. (6) is
@@
which belongs to the PSAHARA family introduced later in Section 3. This is further illustrated in Figure 1.
@@
We can further delve into the risk aversion behaviors of the manager in Eq. (8):
$\text{ARA}\left(X_{T}\right)=\left\{\begin{aligned} &\frac{v\alpha}{\sqrt{\beta^{2}+\left(X_{T}-d\right)^{2}}},&&X_{T}\leqslant B_{T};\\ &\frac{\left(w+v\right)\alpha}{\sqrt{\beta^{2}+\left(X_{T}-wB_{T}-d\right)^{2}}},&&X_{T}>B_{T}.\end{aligned}\right.$ (9)
@@ 3.1 Market Model
We consider a multi-dimensional Black–Scholes model. To solve the explicit solution, all the market parameters are deterministic processes. There are a risk-free asset and risky assets in the market. The risk-free asset has a deterministic return rate and no volatility, while the risky assets have higher expected return rates and positive volatilities. We denote by the filtered probability space $\left(\Omega,\mathcal{F}_{T},\{\mathcal{F}_{t}\}_{0\leqslant t\leqslant T},\mathbb{P}\right)$ the financial market. The filtration $\{\mathcal{F}_{t}\}_{0\leqslant t\leqslant T}$ is the one generated by a $q$-dimensional standard independent Brownian motion $\{\mathbf{W}_{t}\}_{0\leqslant t\leqslant T}=\{\left(W_{1,t},\ldots,W_{q,t}\right)^{\intercal}\}_{0\leqslant t\leqslant T}$ and further augmented by all $\mathbb{P}$-null sets. Let the deterministic process $\{r_{t}\}_{0\leqslant t\leqslant T}$ denote the risk-free rate. The risk-free asset $\{S_{0,t}\}_{0\leqslant t\leqslant T}$ satisfies
@@
In the traditional incomplete Black–Scholes model, the market price of risk $\{\tilde{\boldsymbol{\theta}}_{t}\}_{0\leqslant t\leqslant T}$ can be any solution of the linear system:
$\boldsymbol{\sigma}_{t}\tilde{\boldsymbol{\theta}}_{t}=\boldsymbol{\mu}_{t}-r_{t}\mathbf{1}_{m}.$ (13)
Hence, if $m<q$, $\tilde{\boldsymbol{\theta}}_{t}$ is not unique. We choose Eq. (12), one of the solutions of Eq. (13), to represent the price of risk. Later in Section 4.2, we will show that this simplification does not influence the validity of our main theorems.
We proceed to give a standing assumption on market coefficients. The following is supposed to hold in the rest of the paper.
Assumption 1.
The deterministic processes $\{\boldsymbol{\mu}_{t}\}_{0\leqslant t\leqslant T}$, $\{r_{t}\}_{0\leqslant t\leqslant T}$ and $\{\tilde{\boldsymbol{\theta}}_{t}\}_{0\leqslant t\leqslant T}$ in (13) satisfy
$\sup_{t\in[0,T]}\|\boldsymbol{\mu}_{t}\|_{2}<\infty,\quad\sup_{t\in[0,T]}r_{t}<\infty,\quad\sup_{t\in[0,T]}\|\tilde{\boldsymbol{\theta}}_{t}\|_{2}<\infty,$ (14)
and
$\exp\left\{\frac{1}{2}\int_{0}^{T}\|\boldsymbol{\theta}_{t}\|_{2}^{2}\,\mathrm{d}t\right\}<\infty,$ (15)
where $\|\boldsymbol{\theta}_{t}\|_{2}:=\left(\sum_{i=1}^{q}\theta_{i,t}^{2}\right)^{\frac{1}{2}}$ represents the $L^{2}$ norm of a vector. Moreover, for the matrix–valued process $\{\boldsymbol{\sigma}_{t}\}_{0\leqslant t\leqslant T}$, let $\lambda_{t}^{\max}$ denote the largest eigenvalue of $\boldsymbol{\sigma}_{t}\boldsymbol{\sigma}_{t}^{\intercal}$. Then the process $\{\lambda_{t}^{\max}\}_{0\leqslant t\leqslant T}$ satisfies
$\sup_{t\in[0,T]}\lambda_{t}^{\max}<\infty.$ (16)
The assumption above is reasonable since the coefficients are deterministic processes. Next, we define one pricing kernel process $\{\xi_{t}\}_{0\leqslant t\leqslant T}$ as follows:
$\xi_{t}:=\exp\left\{-\int_{0}^{t}\left(r_{s}+\frac{1}{2}\|\boldsymbol{\theta}_{s}\|_{2}^{2}\right)\,\mathrm{d}s-\int_{0}^{t}\boldsymbol{\theta}_{s}^{\intercal}\,\mathrm{d}\mathbf{W}_{s}\right\},\quad 0\leqslant t\leqslant T.$ (17)
We denote by $\{\pi_{i,t}\}_{0\leqslant t\leqslant T}$ the amount of money invested in $i$th risky asset $S_{i}$ at time $t$. The wealth process $\{X_{t}\}_{0\leqslant t\leqslant T}$ is uniquely determined by the investment process $\{\boldsymbol{\pi}_{t}=\left(\pi_{1,t},\ldots,\pi_{m,t}\right)\}_{0\leqslant t\leqslant T}$ and an initial value $x_{0}\in\mathbb{R}$:
@@
To define the admissible set of controls, we introduce the following assumptions on $\{\boldsymbol{\pi}_{t}\}_{0\leqslant t\leqslant T}$. We will suppose either of Assumptions 2 or 3 holds in the rest of the paper.
Assumption 2.
There exists a uniform $\varepsilon>0$ such that
$\mathbb{E}\left[\int_{0}^{T}\|\boldsymbol{\pi}_{t}\|_{2}^{2+\varepsilon}\,\mathrm{d}t\right]<\infty\quad\text{for all}\;\{\boldsymbol{\pi}_{t}\}_{0\leqslant t\leqslant T}.$ (19)
Assumption 3.
There exists a uniform constant $C\in\mathbb{R}$ such that the wealth process $\{X_{t}\}_{0\leqslant t\leqslant T}$ controlled by any $\{\boldsymbol{\pi}_{t}\}_{0\leqslant t\leqslant T}$ satisfies
$X_{t}\geqslant C\;\text{a.s.},\quad 0\leqslant t\leqslant T.$ (20)
Remark 5.
As we will show later in Section 4.2, either of the two assumptions guarantees the validity of martingale and duality method. Assumption 3 is widely adopted in current studies and reflected in utility functions; see, e.g., Karatzas, Lehoczky and Shreve (1987), where $C$ is taken as 00. We present Assumption 2 to show a rather explicit condition on $\{\boldsymbol{\pi}_{t}\}_{0\leqslant t\leqslant T}$ for the martingale and duality method to work.
We will discuss later in Section 4.2 about the details of these two assumptions, respectively. Now, we define the admissible set $\mathcal{V}$ of controls.
$\mathcal{V}:=\left\{\boldsymbol{\pi}:[0,T]\times\Omega\rightarrow\mathbb{R}^{m}\right.|$ $\boldsymbol{\pi}\;\text{is}\;\{\mathcal{F}_{t}\}_{0\leqslant t\leqslant T}-\text{progressively measurable and satisfies}$ (21)
[…]
A wealth process is called admissible if it is controlled by an admissible control. The decision maker conducts portfolio selection by solving the expected utility maximization problem:
@@
where $U$ is the composed utility function.